OpenStudy (anonymous):

linear congruence: complete soln: 4x(is equivalent to) 5(mod 21) linear congruence: complete soln: 4x(is equivalent to) 5(mod 21) @Mathematics

OpenStudy (anonymous):

by eea i got x = 5x-4n

OpenStudy (anonymous):

now what?

OpenStudy (amistre64):

4x = 5 mod21 4x = 21x + 5 i believe, unless im reading it wrong

OpenStudy (anonymous):

no 4x+21y=5

OpenStudy (amistre64):

i have to try to recall the mods :)

OpenStudy (anonymous):

i agree mwgc :)

OpenStudy (amistre64):

that looks good

OpenStudy (anonymous):

but my teacher said somethin bout odds and evens

OpenStudy (amistre64):

p(x) = q(x) + r(x) p(x) - q(x) = r(x) ... or some such right

OpenStudy (anonymous):

like 2m+1 and 2m

OpenStudy (amistre64):

its not ringing a bell ..... sorry

OpenStudy (anonymous):

k thx

OpenStudy (anonymous):

ill post again